OpenStudy (anonymous):

someone please help!!! ill give a medal to whoever answers this question. The product of (a − b)(a − b) is a perfect square trinomial. Sometimes Always Never

OpenStudy (mertsj):

A perfect square trinomial is one whose factors are (a-b)^2 or (a+b)^2

OpenStudy (anonymous):

a perfect square trinomial like I said before is \[(a+b)^2=a^2+2ab+b^2\]

OpenStudy (anonymous):

so it is always or sometimes?

OpenStudy (anonymous):

or \[(a-b)^2=a^2-2ab+b^2\] if you mutltiply (a-b)(a-b) an you get one of these two, then yea its always